### v4

@@ 5.2 Jailbreak Attacks
TABLE VII: A summary of attacks and defenses for VLMs.
@@
IDEATOR [253] 2024 Black-box Red teaming+VL LLaVA/InstructBLIP/MiniGPT-4 AdvBench/VAJM test set
@@ 5.2.2 Black-box Attacks
Manually-designed Attacks can be as effective as optimized ones. For instance, FigStep [250] introduces an algorithm that bypasses safety measures by converting harmful text into images via typography, enabling VLMs to visually interpret the harmful intent. VRP [252] adopts a visual role-play approach, using LLM-generated images of high-risk characters based on detailed descriptions. By pairing these images with benign role-play instructions, VRP exploits the negative traits of the characters to deceive VLMs into generating harmful outputs.
@@ 5.3.2 Jailbreak Prevention
AdaShield [256] defends against structure-based jailbreaks by prepending defense prompts to inputs, refining them adaptively through collaboration between the VLM and an LLM-based prompt generator, without requiring fine-tuning. ECSO [258] offers a training-free protection by converting unsafe images into text descriptions, activating the safety alignment of pre-trained LLMs within VLMs to ensure safer outputs. InferAligner [259] applies cross-model guidance during inference, adjusting activations using safety vectors to generate safe and reliable outputs. BlueSuffix [260] introduces a reinforcement learning-based black-box defense framework consisting of three key components: (1) an image purifier for securing visual inputs, (2) a text purifier for safeguarding textual inputs, and (3) a reinforcement fine-tuning-based suffix generator that leverages bimodal gradients to enhance cross-modal robustness.
@@ 5.7 Datasets & Benchmarks
TABLE VIII: Safety and robustness benchmarks for VLMs.
@@
The datasets used in VLM safety research are detailed in Table 2. Below, we review the benchmarks proposed for evaluating VLM safety and robustness, summarized in Table VIII. SafeSight [476] introduces two VQA datasets, OODCV-VQA and Sketchy-VQA, to evaluate out-of-distribution (OOD) robustness, highlighting VLMs’ vulnerabilities to OOD texts and vision encoder weaknesses. MM-SafetyBench [477] focuses on image-based manipulations, revealing vulnerabilities in multi-modal interactions. AVIBench [478] evaluates VLM robustness against 260K adversarial visual instructions, exposing susceptibility to image-based, text-based, and content-biased adversarial visual instructions (AVIs). Jailbreak Evaluation of GPT-4o [479] tests GPT-4o with multi-modal and unimodal jailbreak attacks, uncovering alignment vulnerabilities. JailBreakV-28K [438] assesses the transferability of LLM jailbreak techniques to VLMs, showing high attack success rates across 10 open-source models. These studies collectively reveal significant vulnerabilities in VLMs to OOD inputs, adversarial instructions, and multi-modal jailbreaks.
@@ 6.1 Adversarial Attacks
TABLE IX: A summary of attacks and defenses for Diffusion Models (Part I).
@@
JPA[284] 2024 Grey-box Target Internal Defenses Stable Diffusion, Midjourney, DALL·E 2, PIXART-$\alpha$ I2P
@@
Jailbreak Defense SDIDLD[316] 2024 Inference Guidance Latent space Stable Diffusion MS COCO, I2P, CelebA, Winobias, self-constructed
@@
Diff-Cleanse[334] 2024 Removal DDPM, DDIM, LDM MNIST, CIFAR-10, CelebA-HQ
@@ 6.1.1 White-box Attacks
White-box attacks on T2I diffusion models assume full access to model parameters, allowing direct optimization of text prompts or latent space to degrade or disrupt image generation. For example, SAGE [270] explores both the discrete prompt and latent spaces to uncover failure modes in T2I models, including distorted generations and targeted manipulations. ATM [269] generates attack prompts similar to clean prompts by replacing or extending words using Gumbel Softmax, preventing the model from generating desired subjects.
@@ 6.1.2 Gray-box Attacks
QFA [271] minimizes cosine similarity between original and perturbed text embeddings to generate images that differ as much as possible from the original text. RVTA [272] maximizes image-text similarity to align adversarial prompts with reference images generated by a surrogate diffusion model. MMP-Attack [487] simultaneously maximizes the cosine similarity between the perturbed text embedding and the target embedding in both the text and image modalities, while employing a straight-through estimator to execute the optimization process.
@@ 6.2.3 Black-box Attacks
External Safety Attacks, in the black-box setting, use hand-crafted or LLM-assisted adversarial prompts to mislead the victim model to generate NSFW content. UD [287] highlights the risk of T2I models generating unsafe content, especially hateful memes, by refining unsafe prompts manually. SneakyPrompt [286] uses reinforcement learning to optimize adversarial prompts, which updates its policy network based on filter evasion and semantic alignment. Other methods employ LLMs to refine adversarial prompts. Groot [289] decomposes prompts into objects and attributes to dilute sensitive content. DACA [290] breaks down and recombines prompts using LLMs. SurrogatePrompt [291] targets Midjourney, substituting sensitive terms and leveraging image-to-text modules to generate harmful content at scale. Atlas [288] automates the attack with a two-agent system: one VLM generates adversarial prompts, while an LLM evaluates and selects the best candidates. These LLM-assisted strategies can significantly improve the effectiveness and stealthiness of the attacks.
@@ 6.3.1 Concept Erasure
Anchor-based Erasing is a targeted approach that guides the model to shift the target (undesirable concept) towards a good concept (anchor) by aligning predicted latent noise. AC [295] defines anchor concepts as broader categories encompassing the target concepts (e.g., “Grumpy Cat”“Cat”) and uses standard diffusion loss on text-image pairs of anchors to preserve their integrity while erasing target concepts. ABO [296] removes specific target concepts by modifying classifier guidance, using both explicit (replacing the target with a predefined substitute) and implicit (suppressing attention maps) erasing signals, and includes a penalty term to maintain generation quality. DoCo [297] improves generalization by aligning target and anchor concepts through adversarial training and mitigating gradient conflicts with concept-preserving gradient surgery. SPM [293] uses a 1D adapter and negative guidance [292] to suppress target concepts while ensuring non-target concepts remain consistent, affecting only relevant synonyms. SA [298] applies generative replay and elastic weight consolidation to stabilize model weights and maintain normal generation capabilities while preserving non-target concepts.
@@
Adversarial Erasing enhances previous methods by introducing perturbations to the target concept’s text embedding and using adversarial training to improve robustness. Receler [299] employs a lightweight eraser and adversarial prompt embeddings, iteratively training against each other, while applying a binary mask from U-Net attention maps to target only the concept regions. AdvUnlearn [301] shifts adversarial attacks to the text encoder, targeting the embedding space and using regularization to preserve normal generation. RACE [300] improves efficiency by conducting adversarial attacks at a single timestep, reducing computational complexity. These methods enhance the model’s resistance to adversarial prompts aimed at regenerating erased concepts.
@@ 6.3.2 Inference Guidance
This type of guidance use additional input text to steer the model toward safe content. SLD [314] adjusts noise predictions during inference based on a text condition and unsafe concepts, guiding generation towards the intended prompt while avoiding unsafe content, without requiring fine-tuning. It also introduces the I2P benchmark, a dataset for testing inappropriate content generation.
@@
This approach uses additional implicit representations in the latent space to guide generation. SDIDLD [316] employs self-supervised learning to identify the opposite latent direction of inappropriate concepts (e.g., "anti-sexual") and adds these vectors at the bottleneck layer, preventing harmful content generation.
@@ 6.5.1 Backdoor Detection
Most early research focuses on detecting or validating backdoor triggers. T2IShield [330] is the first backdoor detection and mitigation framework for diffusion models, leveraging the assimilation phenomenon in cross-attention maps, where a trigger suppresses other tokens to generate specific content. Ufid [331] validates triggers by noting that clean generations are sensitive to small perturbations, while backdoor-triggered outputs are more robust. DisDet [332] proposes a low-cost detection method that distinguishes poisoned input noise from clean Gaussian noise by identifying distribution shifts.

### v5

@@ 5.2 Jailbreak Attacks
TABLE VIII: A summary of attacks and defenses for VLMs.
@@
DynVLA [633] 2025 Black-box Transfer-based+VL InstructBLIP/MiniGPT4/LLaVA/Gemini MS-COCO/VQA-v2
@@
CAVALRY-V [634] 2025 Black-box Generator-based+V GPT-4.1/Gemini/QwenVL/InternVL/LLaVA/Aria/MiniCPM MMBench-Video/Video-MME
@@
HIMRD [635] 2024 Black-box Manual pipeline+VL LLaVA/DeepSeek/GPT-4o/Gemini/Qwen-VL SafeBench/tiny-SafeBench
IDEATOR [284] 2025 Black-box Red teaming+VL LLaVA/InstructBLIP/MiniGPT-4 AdvBench/VAJM test set
@@
DPS [636] 2025 Prevention Prevention+V Qwen-VL-Plus/GPT-4o/Gemini-1.5-Flash RTA-100/MultiTrust/Self-Gen/MM-SafetyBench/HADES/VisualAttack
ETA [637] 2025 Prevention Prevention+VL LLaVA/InternVL/InternLM-XComposer/LLaMA3.2-Vision SPA-VL/MM-SafetyBench/FigStep
@@ 5.2.2 Black-box Attacks
Manually-designed Attacks can be as effective as optimized ones. For instance, FigStep [281] introduces an algorithm that bypasses safety measures by converting harmful text into images via typography, enabling VLMs to visually interpret the harmful intent. VRP [283] adopts a visual role-play approach, using LLM-generated images of high-risk characters based on detailed descriptions. By pairing these images with benign role-play instructions, VRP exploits the negative traits of the characters to deceive VLMs into generating harmful outputs. HIMRD [635] introduces a heuristic-induced multimodal risk distribution framework that decomposes harmful prompts into semantically benign text and image components. A two-stage heuristic search then guides the model to reconstruct and affirm the underlying malicious intent.
@@ 5.3.2 Jailbreak Prevention
AdaShield [287] defends against structure-based jailbreaks by prepending defense prompts to inputs, refining them adaptively through collaboration between the VLM and an LLM-based prompt generator, without requiring fine-tuning. ECSO [289] offers a training-free protection by converting unsafe images into text descriptions, activating the safety alignment of pre-trained LLMs within VLMs to ensure safer outputs. InferAligner [290] applies cross-model guidance during inference, adjusting activations using safety vectors to generate safe and reliable outputs. BlueSuffix [291] introduces a reinforcement learning-based black-box defense framework consisting of three key components: (1) an image purifier for securing visual inputs, (2) a text purifier for safeguarding textual inputs, and (3) a reinforcement fine-tuning-based suffix generator that leverages bimodal gradients to enhance cross-modal robustness. DPS [636] introduces a black-box, training-free defense that supervises VLMs using responses from partially cropped images, boosting robustness to visual jailbreaks while maintaining benign performance. ETA [637] presents a two-stage inference-time alignment framework: it first screens the safety of inputs and outputs, then applies alignment through interference prefixes and best-of-N sentence search, enabling safe and helpful responses without extra training.
@@ 5.7 Datasets & Benchmarks
TABLE IX: Safety and robustness benchmarks for VLMs.
@@
MIS [643] 2025 6,185 14
VLJailbreakBench [284] 2025 3,654 11
Argus Inspection [644] 2025 1,430 26
The datasets used in VLM safety research are detailed in Table 2. Below, we review the benchmarks proposed for evaluating VLM safety and robustness, summarized in Table IX. SafeSight [639] introduces two VQA datasets, OODCV-VQA and Sketchy-VQA, to evaluate out-of-distribution (OOD) robustness, highlighting VLMs’ vulnerabilities to OOD texts and vision encoder weaknesses. MM-SafetyBench [640] focuses on image-based manipulations, revealing vulnerabilities in multi-modal interactions. AVIBench [641] evaluates VLM robustness against 260K adversarial visual instructions, exposing susceptibility to image-based, text-based, and content-biased adversarial visual instructions (AVIs). Jailbreak Evaluation of GPT-4o [642] tests GPT-4o with multi-modal and unimodal jailbreak attacks, uncovering alignment vulnerabilities. JailBreakV-28K [594] assesses the transferability of LLM jailbreak techniques to VLMs, showing high attack success rates across 10 open-source models. These studies collectively reveal significant vulnerabilities in VLMs to OOD inputs, adversarial instructions, and multi-modal jailbreaks. MIS [643] presents the first multi-image safety dataset for assessing VLMs’ visual reasoning in complex unsafe scenarios, exposing safety reasoning gaps in current fine-tuning methods and providing nuanced multi-image benchmarks. VLJailbreakBench [284] offers 3,654 adversarial image-text pairs generated by the IDEATOR red-teaming framework to evaluate VLMs under black-box multimodal jailbreak settings. Argus Inspection [644] introduces a detail-oriented benchmark for commonsense safety, embedding causally critical yet textually implicit visual “traps” within realistic scenarios.
@@ 6.1 Adversarial Attacks
TABLE X: A summary of attacks and defenses for Diffusion Models (Part I).
@@
MMP-Attack[305] 2025 Grey-box Similarity-driven Stable Diffusion, DALL-E 3, Imagine Art MS COCO
DORMANT[306] 2025 Grey-box Distance-driven Animate Anyone, MagicAnimate, MagicPose, MusePose, Champ, MuseV, UniAnimate, and ControlNeXt TikTok, Champ, UBC Fashion, and TED Talks
@@
FOOLSDEDIT[302] 2024 White-box Classifier-driven SDEdit CelebAMask-HQ, FFHQ
@@
PGJ[326] 2025 Black-box Target External Defenses Stable Diffusion, DALL-E 2, DALL-E 3, Cogview3, Tongyiwanxiang, Hunyuan self-constructed
R2A[327] 2025 Black-box Target External Defense Stable Diffusion, FLUX, DALL·E 3, Midjourney self-constructed
[…]
@@
SafeGen [343] 2024 Concept Erasure Fine-tuning Stable Diffusion MS COCO, I2P, SneakyPrompt-Dataset, NSFW-56k
CPE [344] 2025 Concept Erasure Fine-tuning Stable Diffusion MS COCO, I2P, MACE-Dataset
@@
PromptGuard[353] 2025 Inference Guidance Input Stable Diffusion MS COCO, I2P, SneakyPrompt-Dataset
@@
SDIDLD[355] 2024 Inference Guidance Latent space Stable Diffusion MS COCO, I2P, CelebA, Winobias, self-constructed
Jailbreak Defense CC[356] 2025 Inference Guidance Latent space Stable Diffusion MS COCO, I2P, self-constructed
@@
Diff-Cleanse[374] 2024 Removal Detect & Remove DDPM, DDIM, LDM MNIST, CIFAR-10, CelebA-HQ
@@
NaviDet[377] 2025 Removal Detect & Remove Stable Diffusion MS-COCO
@@ 6.1.1 White-box Attacks
White-box attacks on T2I diffusion models assume full access to model parameters, allowing direct optimization of text prompts or latent space to degrade or disrupt image generation. For example, SAGE [301] explores both the discrete prompt and latent spaces to uncover failure modes in T2I models, including distorted generations and targeted manipulations. ATM [300] generates attack prompts similar to clean prompts by replacing or extending words using Gumbel Softmax, preventing the model from generating desired subjects. FOOLSDEDIT [302] imperceptibly modifies stroke images by applying a mix of four operations: exposure, motion blur, identity mapping, and an empty operation. It automatically selects the optimal combination to steer SDEdit’s [652] outputs toward a desired attribute, while ensuring that the strokes appear visually unchanged.
@@ 6.1.2 Gray-box Attacks
QFA [303] minimizes cosine similarity between original and perturbed text embeddings to generate images that differ as much as possible from the original text. RVTA [304] maximizes image-text similarity to align adversarial prompts with reference images generated by a surrogate diffusion model. MMP-Attack [305] simultaneously maximizes the cosine similarity between the perturbed text embedding and the target embedding in both the text and image modalities, while employing a straight-through estimator to execute the optimization process. DORMANT  [306] embeds imperceptible PGD noise, optimized with VAE-latent, CLIP-semantic, ReferenceNet-detail, and frame-consistency losses. This causes pose-driven portrait animation models to generate identity-shifted and jittery videos, while the source photo remains visually unchanged.
@@ 6.2.3 Black-box Attacks
External Safety Attacks, in the black-box setting, use hand-crafted or LLM-assisted adversarial prompts to mislead the victim model to generate NSFW content. UD [321] highlights the risk of T2I models generating unsafe content, especially hateful memes, by refining unsafe prompts manually. SneakyPrompt [320] uses reinforcement learning to optimize adversarial prompts, which updates its policy network based on filter evasion and semantic alignment. Other methods employ LLMs to refine adversarial prompts. Groot [323] decomposes prompts into objects and attributes to dilute sensitive content. DACA [324] breaks down and recombines prompts using LLMs. SurrogatePrompt [325] targets Midjourney, substituting sensitive terms and leveraging image-to-text modules to generate harmful content at scale. Atlas [322] automates the attack with a two-agent system: one VLM generates adversarial prompts, while an LLM evaluates and selects the best candidates. These LLM-assisted strategies can significantly improve the effectiveness and stealthiness of the attacks. PGJ [326] identifies unsafe tokens and replaces them with perceptually similar but semantically distant phrases, producing short, natural prompts that evade text filters without directly querying the T2I model. R2A [327] further improves an LLM’s reasoning for jailbreaking T2I models by first fine-tuning on Chain-of-Thought examples based on contextual word meanings, and then applying reinforcement learning guided by a dense attack process reward.
@@ 6.3.1 Concept Erasure
Anchor-based Erasing is a targeted approach that guides the model to shift the target (undesirable concept) towards a good concept (anchor) by aligning predicted latent noise. AC [331] defines anchor concepts as broader categories encompassing the target concepts (e.g., “Grumpy Cat”“Cat”) and uses standard diffusion loss on text-image pairs of anchors to preserve their integrity while erasing target concepts. ABO [332] removes specific target concepts by modifying classifier guidance, using both explicit (replacing the target with a predefined substitute) and implicit (suppressing attention maps) erasing signals, and includes a penalty term to maintain generation quality. DoCo [333] improves generalization by aligning target and anchor concepts through adversarial training and mitigating gradient conflicts with concept-preserving gradient surgery. SPM [329] uses a 1D adapter and negative guidance [328] to suppress target concepts while ensuring non-target concepts remain consistent, affecting only relevant synonyms. SA [334] applies generative replay and elastic weight consolidation to stabilize model weights and maintain normal generation capabilities while preserving non-target concepts. SafeGen [343] fine-tunes the vision-only self-attention of Stable Diffusion on <nude, mosaic, benign> triplets, encouraging nude features to be transformed into mosaics while preserving benign images.
@@
Adversarial Erasing enhances previous methods by introducing perturbations to the target concept’s text embedding and using adversarial training to improve robustness. Receler [335] employs a lightweight eraser and adversarial prompt embeddings, iteratively training against each other, while applying a binary mask from U-Net attention maps to target only the concept regions. AdvUnlearn [337] shifts adversarial attacks to the text encoder, targeting the embedding space and using regularization to preserve normal generation. RACE [336] improves efficiency by conducting adversarial attacks at a single timestep, reducing computational complexity. These methods enhance the model’s resistance to adversarial prompts aimed at regenerating erased concepts. CPE [344] introduces a Residual Attention Gate (ResAG) that activates exclusively on target-concept tokens to precisely erase them. The gate is further strengthened through adversarial embedding attack–defense iterations, providing robust protection.
@@ 6.3.2 Inference Guidance
This type of guidance use additional input text to steer the model toward safe content. SLD [352] adjusts noise predictions during inference based on a text condition and unsafe concepts, guiding generation towards the intended prompt while avoiding unsafe content, without requiring fine-tuning. It also introduces the I2P benchmark, a dataset for testing inappropriate content generation. PromptGuard [353] learns a safety soft prompt within the text embedding space and appends it as a suffix to every user prompt, steering the T2I model away from NSFW outputs without altering its weights.
@@
This approach uses additional implicit representations in the latent space to guide generation. SDIDLD [355] employs self-supervised learning to identify the opposite latent direction of inappropriate concepts (e.g., "anti-sexual") and adds these vectors at the bottleneck layer, preventing harmful content generation. Concept Corrector [356] functions during image generation by employing a Generation Check Mechanism (GCM) to inspect an intermediate prediction of the final image for unwanted concepts. If such concepts are detected, a Concept Removal Attention (CRA) module is then activated to dynamically replace the target features with those associated with a negative concept.
@@ 6.5.1 Backdoor Detection
Most early research focuses on detecting or validating backdoor triggers. T2IShield [370] is the first backdoor detection and mitigation framework for diffusion models, leveraging the assimilation phenomenon in cross-attention maps, where a trigger suppresses other tokens to generate specific content. Similarly, NaviDet [377] detects trigger samples by identifying unusual activations in the early diffusion steps induced by specific input tokens. Ufid [371] validates triggers by noting that clean generations are sensitive to small perturbations, while backdoor-triggered outputs are more robust. DisDet [372] proposes a low-cost detection method that distinguishes poisoned input noise from clean Gaussian noise by identifying distribution shifts.